i throw a ball into the air in a movie theater.Ithrow it up with an initial velocity of 64 feet from a balcony 10 feet high .The ceiling of the theater is 58 feet high.When will the ball hit the ceiling

Answer by Greenfinch(383)

Distance S is 58 - 10 feet
Acceleration is -32 feet/sec^2
Initial velocity is 64 feet/sec (not feet as in the question)
so using S = ut - (1/2)at^2 gives 48 = 64t - 16t^2
divide by 16 to get t^2 - 4t + 3 = 0
(t - 3)(t - 1)=0
So t = 1 or 3
It therefore hits at t = 1 second ( the other is the time it would hit if the ball were to pierce the ceiling without losing energy and then hit it again on the way down).